# Q2E: Query-to-Event Decomposition for Zero-Shot Multilingual Text-to-Video Retrieval

### v1

@@ A.1.1 Effect of VLM Sizes
In Q2E, we utilized the InternVL-2.5-78B model. Similar to the Llama-3.3-70B, this model is highly robust and generalizable but comes with significant computational and time constraints. To assess the efficiency and flexibility of our method, Table 6 reports results across various VLM sizes within the InternVL-2.5 family.
@@
A.1.2 Effect of Number of Frames
@@
Figure 3: Impact of the number of frames on retrieval performance. The figure shows the nDCG performance against comparisons per second for three methods: MultiCLIP (black), Q2E (w/o ASR) (blue), and the Q2E (w ASR) (orange). The point size is proportional to the number of sampled frames (i.e., 2, 4, 8, 16, 32, 64). Comparison per Second = Runtime / (num. of Query $\times$ num. of Video). Results are reported on the MultiVENT dataset. The full table is reported on Table 7.
Fig. 3 illustrates the impact of the number of frames on performance across different methods. The results show that increasing the number of sampled frames generally improves performance. However, as the number of comparisons per second increases (i.e., fewer frames are sampled), NDCG declines across all methods, highlighting a trade-off between computational efficiency and retrieval performance.
@@
A.1.3 Effect of Frame Selection Methods
@@
In Q2E, we employed a uniform sampling method for frame selection. However, to compare its effectiveness against a more complex approach, we conducted an ablation study using an adaptive detector from PySceneDetector [^7] https://www.scenedetect.com/. As shown in Table 8, uniform sampling achieves the best performance across both variants, demonstrating its robustness regardless of audio availability. Furthermore, the scene-detection algorithm yields a variable number of frames (ranging from 2 to 100), significantly increasing retrieval time by approximately 2.8 times compared to 16-uniform sampling.
A.1.4 Effect of Different Languages
[…]
Arabic
MultiCLIP + Q2E 10.82 45.86 74.46 90.20 78.43 63.92 0.94 78.09 86.08 14.91 6.0
MultiCLIP + Q2E + ASR 10.79 49.66 79.41 90.20 83.53 68.24 0.94 82.07 88.32 11.59 6.0
Chinese
MultiCLIP + Q2E 10.66 48.81 82.13 98.08 90.00 76.35 0.99 86.32 94.64 9.59 6.0
MultiCLIP + Q2E + ASR 10.46 50.55 83.02 96.15 93.08 76.92 0.97 86.61 93.96 8.73 5.5
English
MultiCLIP + Q2E 10.61 51.66 86.45 100.00 97.69 82.50 1.00 90.57 98.04 8.07 5.0
MultiCLIP + Q2E + ASR 10.61 51.66 87.80 100.00 97.69 83.85 1.00 91.43 98.12 7.74 5.0
Korean
MultiCLIP + Q2E 9.68 43.65 72.65 90.38 82.31 69.04 0.93 76.92 87.04 15.87 6.0
MultiCLIP + Q2E + ASR 9.68 45.93 76.90 90.38 86.15 73.27 0.92 80.58 89.02 13.85 6.0
Russian
MultiCLIP + Q2E 10.39 50.55 81.61 92.31 89.62 73.46 0.96 84.62 91.46 10.88 6.0
MultiCLIP + Q2E + ASR 11.21 53.49 85.40 98.08 94.23 76.54 0.99 88.70 95.02 9.13 5.0
The detailed results for the experiment described in § 5.3 are reported on Table 9.
A.1.5 Effect of LLM Sizes
@@
The detailed results for the experiment described in § 5.4 are reported on Table 10.
A.1.6 Effect of different Rank Fusion
@@
The detailed results for the experiment described in § 5.5 are reported on Table 11. These experiments consider five different approaches for aggregating the scores and performing zero-shot rank fusion. In all approaches, we compute a distribution $P_{i}$ across all videos, for each of the $i$ score components, as discussed in § 5.5. The specific rank fusion approaches we consider then vary in how we combine each of the five $P_{i}$ in order to get a single ranking $\hat{S}$ across all videos. Those approaches are:
@@
This is the approach described in the main paper, in § 5.5.
A.1.7 Effect of Different Components
@@
Q2E  - Video 9.28 39.56 58.00 84.17 72.97 53.90 0.89 64.83 82.66 65.97 8
Q2E  - Query 10.08 45.97 74.74 91.12 83.78 68.80 0.93 78.78 87.92 21.00 6
Q2E  - Events 9.95 46.38 74.77 89.96 84.48 68.88 0.93 79.02 88.98 18.35 6
@@
Q2E  - Video 10.16 44.61 67.74 91.51 81.16 62.43 0.93 73.96 88.97 50.96 6
Q2E  - Query 10.08 47.39 78.12 91.12 85.95 71.74 0.93 81.54 89.18 16.32 6
Q2E  - Events 10.22 48.09 77.77 91.89 87.18 71.47 0.94 81.75 90.38 16.64 6
Table 12: Full analysis of different components in our method. The top block shows the performance of our method with audio transcripts, while the bottom block represents the performance without audio transcripts. The first row in each block corresponds to our full method with all components included. In subsequent rows, one main component (query, video, event) is removed at a time to demonstrate its contribution to the overall method. Events = (Prequel + Current + Sequel). Results are reported on the MultiVENT dataset.
The detailed results for the experiment described in § 5.6 are reported on Table 12.
A.2 Dataset Statistics
The detailed dataset statistics are presented on Table 13.
@@ A.2 Dataset Statistics
Table 13: Overview of the datasets used in this study. The table summarizes key statistics of the MSR-VTT-1kA and MultiVENT datasets.
A.3 Implementation Details
We used Llama-3888meta-llama/Llama-3.3-70B-Instruct as the LLM, InternVL-2.5999OpenGVLab/InternVL2_5-38B as the VLM, and Whisper-large-v3 101010openai/whisper-large-v3 as the multilingual ASR model for query decomposition, frame captioning and audio transcription. For embedding computations, we employed ColBERT111111hltcoe/plaidx-large for text-to-text similarity and MultiCLIP121212laion/Clip-VIT-H14 for text-to-video similarity score calculation. Query decomposition and video captioning were performed with a temperature of 0.8 and top-p sampling of 0.95. For frame extraction, we applied uniform sampling with 16 frames. Our evaluation framework is based on torchmetrics [^13] https://lightning.ai/docs/torchmetrics. The data extraction prompts were executed using two A100 GPUs, while the text-to-video retrieval was conducted on an RTX 6000 GPU.
A.4 Prompts
A.4.1 Extracting Spatial, Temporal, and Event Data
@@
[…]
[Extract Primary Event] You are given a video search query. Your task is to extract the primary event(s) being referenced or discussed in the query.
Here are some guidelines:
- Identify the main event(s) that the query is about, whether explicitly stated or implied
- If multiple events are mentioned, list all of them
- If no clear event can be identified from the query, output "NOT AVAILABLE" in the EVENTS section
- First, explain your reasoning in 1-2 sentences about how you identified the event(s) from the query
- Structure your response into two sections, "EXPLANATION:" and then "EVENTS:"
- The events section should contain either the identified event(s) in a numbered list, or "NOT AVAILABLE" if no event can be determined
- Don’t include temporal or spatial information in this event, just focus on the primary event(s) being discussed or referenced
Here is the query I’d like you to extract events from: {{query}}
[…]
[Extract Spatial Information] You are given a video search query. Assume the video search query is an ongoing event. Your task is to extract any locations mentioned in the query.
Here are some guidelines:
- Extract any location information mentioned in the query, including specific cities, towns, countries, regions, areas, neighborhoods, or other geographical references
- If no location information is found in the query, output "NOT AVAILABLE" in the LOCATION INFORMATION section
- First, explain your reasoning in 1-2 sentences about what location information appears in the query (or lack thereof) and why you’re extracting it
- Structure your response into two sections, "EXPLANATION:" and then "LOCATION INFORMATION:"
- The location information section should contain either the extracted location information in one line, or "NOT AVAILABLE" if no location information is provided.
- Output in the following format: CITY, COUNTRY, REGION, AREA, NEIGHBORHOOD, etc. (e.g. New York, USA, Europe, Downtown). If one of the fields is not available, you can leave it out.
Here is the query I’d like you to extract location information from: {{query}}
[…]
[Extract Temporal Informaton] You are given a video search query. Assume the video search query is an ongoing event. Your task is to extract any dates, times, or years mentioned in the query.
Here are some guidelines:
- Extract any temporal information mentioned in the query, including specific dates, times of day, years, seasons, or other time-related references
- If no temporal information is found in the query, output "NOT AVAILABLE" in the TEMPORAL INFORMATION section
- First, explain your reasoning in 1-2 sentences about what temporal information appears in the query (or lack thereof) and why you’re extracting it
- Structure your response into two sections, "EXPLANATION:" and then "TEMPORAL INFORMATION:"
- The temporal information section should contain either the extracted date/time information in one line, or "NOT AVAILABLE" if no temporal information is provided.
- Output in the following format: DAY MONTH YEAR, TIME, SEASON, etc. (e.g. 1 January 2022, 3:00 PM, Summer). If one of the fields is not available, you can leave it out.
Here is the query I’d like you to extract temporal information from: {{query}}
A.4.2 Decomposing Query to Prequel, Current, and Sequel Events
[…]
[ Prequel / Current / Sequel ] You are given a video search query. Assume the video search query is an ongoing event. You have to find out what previous events could lead to this given event / decompose the ongoing event in multiple simple events / find out what events can be the outcome of this ongoing event.The extracted events should be concrete enough to be visualized in a video.
Here are some guidelines:
- Extract events that could be visualized in a video.
- Before extracting an event, consider whether the event text could serve as a reasonable caption for a frame of a video associated with the query.
- Give specific events related to the query, use your previous WORLD KNOWLEDGE to extract events that are likely to be seen in a video. For example, if the query is "2019 Christmas Eve Protests in Hong Kong", you might return "Protesters marching in the streets", but not "Protestors marching in snow" because it is not likely to have snow in Hong Kong, though its pretty common in other places.
- First, explain your reasoning in 1-2 sentences about which events it makes sense to extract. Then, provide a numbered list with the events (anywhere from one to five event; use your best judgement to determine how many to extract)
- Structure your response into two sections, ""EXPLANATION:"" and then ""EVENTS:"". The events sections should contain nothing but the events.
Here is the query I’d like you to extract events from: {{query}}
A.4.3 Refining Query
@@
[…]
[ Refine Query ] You will receive a query along with optional time, location, and main event information. Your task is to integrate this information naturally into a refined, human-like search query.
Here are some guidelines:
- Take the original query and incorporate any provided time, location, and event details seamlessly
- If any of these elements (time, location, event) are not provided, work with just the available information
- The refined query should sound natural, as if someone is searching with that query in Google, YouTube, etc.
- The base query information must always be preserved in the refined version
- First, explain your reasoning in 1-2 sentences about which events it makes sense to extract. Then, provide the refine query in one line
- Structure your response into two sections, ""EXPLANATION:"" and then ""REFINED QUERY:"". The Refined Query section should contain nothing but the refined query in one line.
Base Query: {{prequel/Current/sequel}}
Event: {{event}}
Place: {{place}}
Time: {{time}}
A.4.4 Frame Captioner
The following prompt was used to generate captions for each video frame. Unlike other prompts, a VLM was utilized to handle the multimodal inputs effectively. As illustrated in Fig. 2, we implemented two variants: (1) ASR and (2) Non-ASR. The Orange texts were excluded in the prompt for the Non-ASR variant.
[…]
[ Frame Caption ] You are provided with the description of the previous frames, the current frame (image) of a single video, and the original transcript of the audio (ASR) from the entire video. Your task is to describe the main event or activity depicted in the current frame.
# Instructions:
- Use the original ASR to enhance your understanding of the video context and inform your description of the current frame.
- Focus on identifying and detailing the key actions, participants, and context of the event depicted in the current frame.
- Maintain continuity with the previous frame’s description, connecting the current frame to the overall flow of the video.
- Ignore unrelated visual elements like colors or shapes unless they are critical to understanding the event.
- Use any visible text in the image to enhance your interpretation of the frame.
Only output the current frame description, integrating insights from the original ASR and previous frame description. Don’t output any additional commentary or content.
# Original ASR Transcript (Entire Video):
{{ASR}}
# Previous Frame Description:
{{prev_caption}}
# Frame:
{{frame}}
# Current Frame Description:
A.4.5 Full Video Captioner

### v2

@@ A.3.1 Effect of VLM Sizes
In Q2E, we utilized the InternVL-2.5-78B model. Similar to the Llama-3.3-70B, this model is highly robust and generalizable but comes with significant computational and time constraints. To assess the efficiency and flexibility of our method, Table˜8 reports results across various VLM sizes within the InternVL-2.5 family.
@@
A.3.2 Effect of different layers in Translation Module
[…]
Q2E 79.60 73.09 0.95 83.24
    w/o ASR 79.30 72.86 0.95 83.02
    w/o Translator 79.47 73.01 0.95 83.14
    w/o Refiner 79.25 72.82 0.95 82.95
    w/o (ASR + Translator + Refiner) 75.76 69.73 0.95 80.04
Table 9: Analysis of different layers of the translation module. Results are reported on the MultiVENT dataset.
As described in the §˜2.3, in our method we have used a multi-layer ASR system consisting of ASR, an automatic translator and an LLM refiner. To assess the impact of different layers of translation, we have provided ablation studies in Table˜9. In summary, while excluding one translation layer results in a small reduction in NDCG score (0.1-0.5 points), excluding all translation layers leads to a much more significant decline (2-7 points). This suggests strong complementary aspects and a strong aggregate effect.
A.3.3 Effect of Number of Frames
@@
Figure 4: Impact of the number of frames on retrieval performance. The figure shows the nDCG performance against comparisons per second for three methods: MultiCLIP (black), Q2E (without ASR) (blue), and the Q2E (with ASR) (orange). The point size is proportional to the number of sampled frames (i.e., 2, 4, 8, 16, 32, 64). Comparison per Second = Runtime / (num. of Query $\times$ num. of Video). Results are reported on the MultiVENT dataset. The full table is reported on Table˜10.
Fig.˜4 illustrates the impact of the number of frames on performance across different methods. The results show that increasing the number of sampled frames generally improves performance. However, as the number of comparisons per second increases (i.e., fewer frames are sampled), NDCG declines across all methods, highlighting a trade-off between computational efficiency and retrieval performance.
@@
A.3.4 Effect of Frame Selection Methods
@@
In Q2E, we employed a uniform sampling method for frame selection. However, to compare its effectiveness against a more complex approach, we conducted an ablation study using an adaptive detector from PySceneDetector [^6] https://www.scenedetect.com/. As shown in Table˜11, uniform sampling achieves the best performance across both variants, demonstrating its robustness regardless of audio availability. Furthermore, the scene-detection algorithm yields a variable number of frames (ranging from 2 to 100), significantly increasing retrieval time by approximately 2.8 times compared to 16-uniform sampling.
A.3.5 Effect of LLM Sizes
@@
The detailed results for the experiment described in §˜3.4 are reported on Table˜12.
A.3.6 Effect of different Rank Fusion
@@
[…]
Mean over events & Max over captions 77.91 71.58 0.95 81.97
Max over events & Mean over captions 0.46 0.39 0.00 0.44
Max over events & Mean of Top 3 captions 78.90 72.47 0.94 82.54
Max over event & Mean over Top 5 captions 78.91 72.47 0.94 82.50
Mean over Top 3 events & Max over captions 79.11 72.70 0.95 82.91
Mean over Top 3 events & Mean over Top 3 captions 78.96 72.51 0.94 82.62
Q2E 79.25 72.82 0.95 82.95
Table 14: Performance comparison of different aggregation methods across retrieval metrics. Q2E, with global maximum, achieves the best performance across all metrics.
The detailed results for the experiment described in §˜3.5 are reported on Tables˜13 and 14. These experiments consider five different approaches for aggregating the scores and performing zero-shot rank fusion. In all approaches, we compute a distribution $P_{i}$ across all videos, for each of the $i$ score components, as discussed in §˜3.5. The specific rank fusion approaches we consider then vary in how we combine each of the five $P_{i}$ in order to get a single ranking $\hat{S}$ across all videos. Those approaches are:
@@
This is the approach described in the main paper, in §˜3.5.
A.3.7 Effect of Different Components
@@
    w/o Video 9.28 39.56 58.00 84.17 72.97 53.90 0.89 64.83 82.66 65.97 8
    w/o Query 10.08 45.97 74.74 91.12 83.78 68.80 0.93 78.78 87.92 21.00 6
    w/o Event 9.95 46.38 74.77 89.96 84.48 68.88 0.93 79.02 88.98 18.35 6
     w/o Prequel 10.11 46.73 75.73 91.12 85.17 69.73 0.94 79.85 89.13 20.13 6
     w/o Current 10.20 46.79 75.85 91.89 85.33 69.81 0.94 80.02 89.29 20.29 6
     w/o Sequel 10.07 46.66 75.75 91.12 85.02 69.73 0.94 79.85 89.19 20.15 6
@@
    w/o Video 10.16 44.61 67.74 91.51 81.16 62.43 0.93 73.96 88.97 50.96 6
    w/o Query 10.08 47.39 78.12 91.12 85.95 71.74 0.93 81.54 89.18 16.32 6
    w/o Event 10.22 48.09 77.77 91.89 87.18 71.47 0.94 81.75 90.38 16.64 6
     w/o Prequel 10.28 48.86 79.65 92.66 88.49 73.17 0.95 83.25 91.04 15.89 6
     w/o Current 10.32 48.78 79.62 93.05 88.34 73.13 0.95 83.25 91.04 15.86 6
     w/o Sequel 10.32 48.91 79.34 93.05 88.57 72.86 0.95 83.09 91.25 15.81 6
Table 15: Full analysis of different components in our method. The top block shows the performance of our method with audio transcripts, while the bottom block represents the performance without audio transcripts. The first row in each block corresponds to our full method with all components included. In subsequent rows, one main component (video, query, Event) is removed at a time to demonstrate its contribution to the overall method. While Event consists of Prequel, Current and Sequel, on the bottom three rows, we have removed each of them one at a time to demonstrate the contribution of each. Results are reported on the MultiVENT dataset.
The detailed results for the experiment described in §˜3.6 are reported on Table˜15. The fine-grained results for each component of events show a slight reduction in NDCG score (0.1-0.3) if we exclude one of the prequel, sequel, or current generations, if we exclude all, we see a much significant reduction (1-2 points), suggesting strong complementary aspects and a strong aggregate effect.
A.4 Dataset Statistics
The detailed dataset statistics are presented on Table˜16.
@@ A.4 Dataset Statistics
Table 16: Overview of the datasets used in this study. The table summarizes key statistics of the MSR-VTT-1kA (Xu et al., 2016), MSVD (Chen and Dolan, 2011b) and MultiVENT (Sanders et al., 2023) datasets.
A.5 Implementation Details
We used Llama-3777meta-llama/Llama-3.3-70B-Instruct as the LLM, InternVL-2.5888OpenGVLab/InternVL2_5-38B as the VLM, and Whisper-large-v3 999openai/whisper-large-v3 as the multilingual ASR model for query decomposition, frame captioning and audio transcription. For embedding computations, we employed ColBERT101010hltcoe/plaidx-large for text-to-text similarity and MultiCLIP111111laion/Clip-VIT-H14 for text-to-video similarity score calculation. Query decomposition and video captioning were performed with a temperature of 0.8 and top-p sampling of 0.95. For frame extraction, we applied uniform sampling with 16 frames. Our evaluation framework is based on torchmetrics [^12] https://lightning.ai/docs/torchmetrics. The data extraction prompts were executed using two A100 GPUs, while the text-to-video retrieval was conducted on an RTX 6000 GPU.
A.6 Prompts
In the following sections, we provide all the prompts used for the event decomposition (§˜A.6.1, A.6.2 and A.6.3), video decomposition (§˜A.6.4 and A.6.5) and ASR module (§˜A.6.6).
A.6.1 Decomposing Query to Prequel, Current, and Sequel Events
Prequel / Current / Sequel You are given a video search query. Assume the video search query is an ongoing event. You have to find out what previous events could lead to this given event / decompose the ongoing event in multiple simple events / find out what events can be the outcome of this ongoing event.The extracted events should be concrete enough to be visualized in a video. Here are some guidelines: - Extract events that could be visualized in a video. - Before extracting an event, consider whether the event text could serve as a reasonable caption for a frame of a video associated with the query. - Give specific events related to the query, use your previous WORLD KNOWLEDGE to extract events that are likely to be seen in a video. For example, if the query is "2019 Christmas Eve Protests in Hong Kong", you might return "Protesters marching in the streets", but not "Protestors marching in snow" because it is not likely to have snow in Hong Kong, though its pretty common in other places. - First, explain your reasoning in 1-2 sentences about which events it makes sense to extract. Then, provide a numbered list with the events (anywhere from one to five event; use your best judgement to determine how many to extract) - Structure your response into two sections, ""EXPLANATION:"" and then ""EVENTS:"". The events sections should contain nothing but the events. Here is the query I’d like you to extract events from: {{query}}
A.6.2 Extracting Spatial, Temporal, and Event Data
@@
Extract Primary Event You are given a video search query. Your task is to extract the primary event(s) being referenced or discussed in the query. Here are some guidelines: - Identify the main event(s) that the query is about, whether explicitly stated or implied - If multiple events are mentioned, list all of them - If no clear event can be identified from the query, output "NOT AVAILABLE" in the EVENTS section - First, explain your reasoning in 1-2 sentences about how you identified the event(s) from the query - Structure your response into two sections, "EXPLANATION:" and then "EVENTS:" - The events section should contain either the identified event(s) in a numbered list, or "NOT AVAILABLE" if no event can be determined - Don’t include temporal or spatial information in this event, just focus on the primary event(s) being discussed or referenced Here is the query I’d like you to extract events from: {{query}}
Extract Spatial Information You are given a video search query. Assume the video search query is an ongoing event. Your task is to extract any locations mentioned in the query. Here are some guidelines: - Extract any location information mentioned in the query, including specific cities, towns, countries, regions, areas, neighborhoods, or other geographical references - If no location information is found in the query, output "NOT AVAILABLE" in the LOCATION INFORMATION section - First, explain your reasoning in 1-2 sentences about what location information appears in the query (or lack thereof) and why you’re extracting it - Structure your response into two sections, "EXPLANATION:" and then "LOCATION INFORMATION:" - The location information section should contain either the extracted location information in one line, or "NOT AVAILABLE" if no location information is provided. - Output in the following format: CITY, COUNTRY, REGION, AREA, NEIGHBORHOOD, etc. (e.g. New York, USA, Europe, Downtown). If one of the fields is not available, you can leave it out. Here is the query I’d like you to extract location information from: {{query}}
Extract Temporal Informaton You are given a video search query. Assume the video search query is an ongoing event. Your task is to extract any dates, times, or years mentioned in the query. Here are some guidelines: - Extract any temporal information mentioned in the query, including specific dates, times of day, years, seasons, or other time-related references - If no temporal information is found in the query, output "NOT AVAILABLE" in the TEMPORAL INFORMATION section - First, explain your reasoning in 1-2 sentences about what temporal information appears in the query (or lack thereof) and why you’re extracting it - Structure your response into two sections, "EXPLANATION:" and then "TEMPORAL INFORMATION:" - The temporal information section should contain either the extracted date/time information in one line, or "NOT AVAILABLE" if no temporal information is provided. - Output in the following format: DAY MONTH YEAR, TIME, SEASON, etc. (e.g. 1 January 2022, 3:00 PM, Summer). If one of the fields is not available, you can leave it out. Here is the query I’d like you to extract temporal information from: {{query}}
A.6.3 Refining Query
@@
Refine Query You will receive a query along with optional time, location, and main event information. Your task is to integrate this information naturally into a refined, human-like search query. Here are some guidelines: - Take the original query and incorporate any provided time, location, and event details seamlessly - If any of these elements (time, location, event) are not provided, work with just the available information - The refined query should sound natural, as if someone is searching with that query in Google, YouTube, etc. - The base query information must always be preserved in the refined version - First, explain your reasoning in 1-2 sentences about which events it makes sense to extract. Then, provide the refine query in one line - Structure your response into two sections, ""EXPLANATION:"" and then ""REFINED QUERY:"". The Refined Query section should contain nothing but the refined query in one line. Base Query: {{prequel/Current/sequel}} Event: {{event}} Place: {{place}} Time: {{time}}
A.6.4 Frame Captioner
The following prompt was used to generate captions for each video frame. Unlike other prompts, a VLM was utilized to handle the multimodal inputs effectively. As illustrated in Fig.˜2, we implemented two variants: (1) ASR and (2) Non-ASR. The Orange texts were excluded in the prompt for the Non-ASR variant.
Frame Caption You are provided with the description of the previous frames, the current frame (image) of a single video, and the original transcript of the audio (ASR) from the entire video. Your task is to describe the main event or activity depicted in the current frame. # Instructions: - Use the original ASR to enhance your understanding of the video context and inform your description of the current frame. - Focus on identifying and detailing the key actions, participants, and context of the event depicted in the current frame. - Maintain continuity with the previous frame’s description, connecting the current frame to the overall flow of the video. - Ignore unrelated visual elements like colors or shapes unless they are critical to understanding the event. - Use any visible text in the image to enhance your interpretation of the frame. Only output the current frame description, integrating insights from the original ASR and previous frame description. Don’t output any additional commentary or content. # Original ASR Transcript (Entire Video): {{ASR}} # Previous Frame Description: {{prev_caption}} # Frame: {{frame}} # Current Frame Description:
A.6.5 Full Video Captioner